For a free electron with 100 keV kinetic energy, calculate the:

a) electron speed

b) electron momentum

c) de Broglie wavelength of the electron
Answer #1

a) Energy in J = Energy in eV * 1.6 * 10-19 = 100 * 103 * 1.6 * 10-19 = 1.6 * 10-14 J

We have the relation : E = 1/2 * m * v2, where m is mass and v is velocity

We use the above energy and value of m to solve for v:

1.6 * 10-14 = 1/2 * 9.11 * 10-31 * v2

Solving, we get: v = 7.02 * 108 m/s

b) momentum = m * v = 9.11 * 10-31 * 7.02 * 108 = 6.40 * 10-22 kg m/s

c) De broglie waverlength = h / momentum, where h is planck's constant

= 6.625 * 10-34 / (6.40 * 10-22 ) = 1.04 * 10-12 m